OpenStudy (anonymous):

Could you please explain and show working for.... Find the derivative of 1. 3(2xsquared +5)sixth power -cos(x)squared 2. cos((xsquared +1)cubed) Thank you

OpenStudy (anonymous):

1) \[18(x ^{2}+5)*2x+2cosx *sinx\]

OpenStudy (anonymous):

2) \[d/dx (\cos(x ^{2}+1)^{3})=-\sin(x ^{2}+1)^{2}*3(x ^{2}+1)*2x\]

OpenStudy (anonymous):

I'm not sure about 2) ... is it (x^2+1)^3 or cos^3...?

OpenStudy (anonymous):

thank you...but should it be 18(2x squared +5)fifth etc....how do you write in powers here?

OpenStudy (anonymous):

you are right.. I forgot fifth power. I use " ^" a lot of people on this site too :) check the icon "Equation" below this box as well

OpenStudy (anonymous):

by the way... my answer 2) I missed second power in (x^2+1) as well... sorry :)

OpenStudy (anonymous):

all good, thank you, :)